could you help me solve the following ellipse problem? The problem gives me that the eccentricity is 1/4. it also says that the foci are (3,2) and (3,10). I can not even think of the first thing to do to this. Could you help me by tomorrow because it is due tomorrow and also could you send me a step to step solution. Thank you, Alena
Found 2 solutions by fastblue, stanbon:
Answer by fastblue(13)

Hi,
Repost what the question is and will solve for you..

Answer by stanbon(75887)

could you help me solve the following ellipse problem? The problem gives me that the eccentricity is 1/4. it also says that the foci are (3,2) and (3,10).
----------------------------
Plot the focus points on an xy-coordinate system.
Find the center halfway between the foci at (3,6)
So c = distance from center to focus = 4
------------------
c = ea = (1/4)a = 4
So a=16
a^2 = 256
--------------
For an ellipse a^2=b^2+c^2
16^2 = b^2 + 4^2
b^2 = 240
----------------------------------
EQUATION:
(x-3)^2/240 + (y-6)^2/256 = 1
==============================
Cheers,
Stan H.